— Simple Tests — Test using accent macros and Latin-1 chars: täter élève élève Test using german hyphenation patterns: täter élève Test using french hyphenation patterns: täter élève Test using french hyphenation patterns (OT1): täter élève délégué — Major Tests —

— T1: À Á Â Ã Ä Ă Å à á â ã ä ă å Ç Ć Č ç ć č Ď Ğ ğ È É Ê Ẽ Ë Ě è é ê ẽ ë ě Ì Í Î Ĩ Ï İ ì í î ĩ ï i ì í î ĩ ï i̇ Ñ Ń Ň ñ ń ň Ò Ó Ô Õ Ö Ő ò ó ô õ ö ő Ŕ Ř ŕ ř Ś Š Ş ś š ş Ť Ţ ţ Ù Ú Û Ũ Ü Ű Ů ù ú û ũ ü ű ů Ý ý Ź Ž Ż ź ž ż

— OT1: À Á Â Ã Ä Ă Å à á â ã ä ă å Ç Ć Č ç ć č Ď Ğ ğ È É Ê Ẽ Ë Ě è é ê ẽ ë ě Ì Í Î Ĩ Ï İ ì í î ĩ ï i̇ ì í î ĩ ï i̇ Ñ Ń Ň ñ ń ň Ò Ó Ô Õ Ö Ő ò ó ô õ ö ő Ŕ Ř ŕ ř Ś Š Ş ś š ş Ť Ţ ţ Ù Ú Û Ũ Ü Ű Ů ù ú û ũ ü ű ů Ý ý Ź Ž Ż ź ž ż

— LO1: À Á Â Ã Ä Ă Å à á â ã ä ă å Ç Ć Č ç ć č Ď Ğ ğ È É Ê Ẽ Ë Ě è é ê ẽ ë ě Ì Í Î Ĩ Ï İ ì í î ĩ ï i ì í î ĩ ï i Ñ Ń Ň ñ ń ň Ò Ó Ô Õ Ö Ő ò ó ô õ ö ő Ŕ Ř ŕ ř Ś Š Ş ś š ş Ť Ţ ţ Ù Ú Û Ũ Ü Ű Ů ù ú û ũ ü ű ů Ý ý Ź Ž Ż ź ž ż